OpenStudy (anonymous):

A rock is dropped from a sea cliff, and the sound of it striking the ocean is heard 3.2 s later. If the speed of sound is 340 m/s, how high is the cliff?

OpenStudy (anonymous):

The right answer's 46 m. xD

OpenStudy (anonymous):

okay, lets see. say at t = 0, the rock is dropped. it hits the ocean at a later time, say t1 the distance of the ocean from the cliff is given by s = ut1+1/2 gt1^2 u = 0 so s = 1/2 * 9.8 *t1^2 say it takes t2 time for the moment the rock hit the ocean to the time you heard the sound. the sound has to travel the same disntace s to reach you. s = 340 t2 so we have 340 t2 = 4.9t1^2 ---------equation 1 and t1+t2 = 3.2 , or t2 = 3.2-t1 -------- equation 2 substitute the value of t2 into equation 1 and find t1. use that to find s

OpenStudy (anonymous):

the distance is 46.02 m

OpenStudy (anonymous):

Would the quadratic formula be required?

OpenStudy (anonymous):

yes use that

OpenStudy (anonymous):

from s=s=vt2=4.9t1^2 340t2=4.9t1^2 69.39 t2=t1^2 substitute t2=3.2-t1

OpenStudy (anonymous):

69.39 t2=t1^2 69.39 (3.2-t1 )=t1^2

OpenStudy (anonymous):

69.39 (3.2-t1 )=t1^2 222.05-69.39t1 =t1^2

OpenStudy (anonymous):

222.05-69.39t1 =t1^2 t1^2+69.39t1 -222.05=0 now use quadratic eq formula

OpenStudy (anonymous):

t1=3.065sec subst this to s=4.9t1^2

OpenStudy (anonymous):

s=4.9(3.065)^2 =46.03 meter ans

OpenStudy (anonymous):

did you get it nubsauce?

OpenStudy (anonymous):

Yeah, thanks for the help.